OpenStudy (anonymous):

Give a unique example of each: 1. Part 1: An absolute value equation with two solutions 2. Part 2: An absolute value equation with one solution 3. Part 3: An absolute value equation with no solutions

OpenStudy (amistre64):

well, no solutions would imply the absolute value is negative one solution is when it equals 0

OpenStudy (amistre64):

im not sure what a "unique" example would be since there are many examples for none and 2

OpenStudy (anonymous):

alright thanks :)

OpenStudy (amistre64):

yep.. now if it said: "Give an example of each:" then its as simple as |x-3| = 5, has 2 solutions |x-3| = 0, only has 1 solution |x-3| = -3, has no solutions